# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ C.4.5 Proof of Lemma 19
[…]
=$\overset{(i)}{=}~{}$ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\|\mathcal{E}_{S,i}\|_{2}^{2}+\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}+2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S},\mathcal{E}_{S,i}\right\rangle$
@@
[…]
@@
    $\operatorname{CN}_{l}^{S}$ :=¯$:=\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2},$
@@
$\mathcal{B}^{(i,1)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{l}^{S}\geq\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}\right\},$ (49)
$\mathcal{B}^{(i,2)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{i}^{S}\geq\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl{(}|S|-2\sqrt{|S|\zeta^{S}}\bigr{)}\right\},$ (50)
$\mathcal{B}^{(i,3)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{ki}^{S}\geq\frac{1-|G_{k}^{\ast}|}{|G_{k}^{\ast}|^{2}}\bigl{(}|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S}\bigr{)}\right\},$ (51)
$\mathcal{B}^{(i,4)}_{kl,S}$ :=$:=\left\{\operatorname{GSN}_{kl}^{S}\geq-2\sqrt{\frac{2\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\},$ (52)
@@
Here, we use the separation condition to bound the term. If we condition on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$ and ¯$\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}$, then $-\mathrm{IP}_{kli}^{S}$ has a univariate Gaussian distribution with mean zero and variance as the following:
@@
=$\overset{(i)}{=}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+\underbrace{\left\|\overline{\mathcal{E}}_{S,l}^{\ast}-(1-|G_{k}^{\ast}|^{-1})^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right\|_{2}^{2}}_{:=\operatorname{CN}_{kli}^{S}}$
@@
$\mathcal{B}^{(i,5)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{kli}^{S}\leq(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})\right\},$
$\mathcal{B}^{(i,6)}_{kl,S}$ :=$:=\left\{\operatorname{GN}_{kli}^{S}\leq 2\sqrt{2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})\zeta^{S}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\}.$
@@
$D^{S}_{kli}(\mathbf{X})~{}\gtrsim$ $~{}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}+\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl{(}|S|-2\sqrt{|S|\zeta^{S}}\bigr{)}$
@@
[…]
@@
[…]
@@
uniformly over $S\in\mathcal{S}$, with probability at least $1-C/n$, where $C>0$ is a constant. This completes the proof of Lemma 19.
@@ C.4.6 Proof of Lemma 20
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}\leq\sup_{\mathbf{v}\in\mathbb{R}^{n},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}=\|\mathcal{E}_{S,\cdot}\|_{\mathrm{op}}^{2}.$
@@ C.4.7 Proof of Lemma 21
$\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}D^{S}_{kli}(\mathbf{X})\gtrsim\underbrace{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|}_{\text{Part A}}\underbrace{-\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}R^{S}_{kl}}_{\text{Part B}},$
@@
$\min_{1\leq k\neq l\leq K}\left\{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}D^{S}_{kli}(\mathbf{X})\right\}$ $\gtrsim\min_{1\leq k\neq l\leq K}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|$
@@ C.4.8 Proof of Lemma 22
$t^{S}_{lk}(\dot{\lambda}^{S})$ =$\overset{(i)}{=}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\dot{\lambda}^{S}\biggr{\}}$
=$\overset{(ii)}{=}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\bigl{(}|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}\bigr{)}\biggr{\}}$
@@
where step $(i)$ follows from the definition of $t^{S}_{lk}(\lambda^{S})$ in (26), step $(ii)$ follows from the definition of $\dot{\lambda}_{S}$ in Lemma 21, step $(iii)$ uses $-\Delta_{S\cap S_{0}}^{2}>-\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}$, step $(iv)$ uses
$-m\geq-2\left\{\frac{|G_{k}^{\ast}|\cdot|G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\},$
@@ C.4.9 Proof of Lemma 23
[…]
@@
$\mathbb{P}\bigl{(}\bigcup_{1\leq k\neq l\leq K}\bigcup_{S\in\mathcal{S}}(\mathcal{C}_{Ikl}^{S})^{C}\bigr{)}\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{Ikl}^{S})^{C})\overset{(i)}{\leq}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\frac{1}{n|S|^{2}{\binom{p}{|S|}}}=\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\frac{1}{nq^{2}}\leq\frac{\pi^{2}}{6n},$
@@ C.4.11 Proof of Lemma 25
$G^{S}_{1,kl}(\mathbf{v})$ $=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$ (59)
$=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\frac{1}{2}(v_{i}+v_{j})\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$
@@
$\mathbb{E}\bigl{[}\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|},\mathbf{X})\bigr{]}=\mathbb{E}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\mathbb{E}\bigl{[}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\bigr{]}=\sum_{i\in G_{k}^{\ast}}v_{i}\mathbb{E}\bigl{[}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{]}\overset{(i)}{=}|S|\sum_{i\in G_{k}^{\ast}}v_{i}=0,$
@@
We start with $\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}$. First note that for $\mathbf{M}\in\mathcal{A}^{S}_{kl}$, its operator norm can be simplified into:
@@
$=\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\sum_{\{(i,j)\in G_{k}^{\ast}\}}u_{i}u_{j}\frac{v_{i}+v_{j}}{2}$
@@
$\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}},\mathbf{X})=\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}+\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}.$
@@
$\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}$ $=\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}\bigr{|}$
@@
$\mathbb{E}\biggl{[}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}\biggr{]}$ $\leq\mathbb{E}\biggl{[}\bigl{(}\sum_{j\in G_{k}^{\ast}}\bigl{\langle}\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr{\rangle}^{2}\bigr{)}^{1/2}\biggr{]}$
@@
=$\stackrel{{\scriptstyle(ii)}}{{=}}\bigl{(}\sum_{j\in G_{k}^{\ast}}\mathbb{E}\|\hskip-6.99997pt\sum_{i\in G_{k}^{\ast},i\neq j}\hskip-6.99997pt\mathcal{E}_{S,i}\|_{2}^{2}\;\bigr{)}^{1/2}$
@@
$\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{|}$ =$\overset{(i)}{=}\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{i\in G_{k}^{*}}v_{i}\bigr{(}\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\bigl{)}\bigr{|}$
@@
$\mathcal{C}_{N,kl}^{S}:=\biggl{\{}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\lesssim\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\underbrace{|G_{k}^{\ast}|\sqrt{|S|}}_{\mathbb{E}\bigl{[}\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\bigr{]}\lesssim}~{}+~{}\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\zeta^{S}_{N,kl}\biggr{\}},$
@@
$\mathbb{P}\biggl{(}(\mathcal{C}_{N,kl}^{S})^{C}\biggr{)}$ $\lesssim\exp\biggl{[}-\min\biggl{(}\underbrace{\frac{(\zeta^{S}_{N,kl})^{2}}{|G_{k}^{\ast}|^{2}|S|}}_{\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{\mathcal{A}^{S}_{kl}}^{2}\lesssim},\quad\underbrace{\frac{\zeta^{S}_{N,kl}}{\sqrt{|G_{k}^{\ast}|}}}_{\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}\lesssim}\biggr{)}\biggr{]}$

### v2

@@ C.4.5 Proof of Lemma 19
$~=$~ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}+\mathcal{E}_{S,i}-\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}-\|\mathcal{E}_{S,i}-\overline{\mathcal{E}}_{S,k}^{\ast}\|_{2}^{2}$
$~=$~ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}+\mathcal{E}_{S,i}-\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}-\left(\frac{|G_{k}^{\ast}|-1}{|G_{k}^{\ast}|}\right)^{2}\|\mathcal{E}_{S,i}-\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\|_{2}^{2}$
$\overset{(i)}{=}$~ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\|\mathcal{E}_{S,i}\|_{2}^{2}+\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}+2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S},\mathcal{E}_{S,i}\right\rangle$
@@
$~=$~ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}+\frac{2|G_{k}^{\ast}|-1}{|G_{k}^{\ast}|^{2}}\|\mathcal{E}_{S,i}\|_{2}^{2}$
@@
    $\operatorname{CN}_{l}^{S}$ $:=\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2},$
@@
$\mathcal{B}^{(i,1)}_{kl,S}$ $:=\left\{\operatorname{CN}_{l}^{S}\geq\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}\right\},$ (53)
$\mathcal{B}^{(i,2)}_{kl,S}$ $:=\left\{\operatorname{CN}_{i}^{S}\geq\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl(|S|-2\sqrt{|S|\zeta^{S}}\bigr)\right\},$ (54)
$\mathcal{B}^{(i,3)}_{kl,S}$ $:=\left\{\operatorname{CN}_{ki}^{S}\geq\frac{1-|G_{k}^{\ast}|}{|G_{k}^{\ast}|^{2}}\bigl(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S}\bigr)\right\},$ (55)
$\mathcal{B}^{(i,4)}_{kl,S}$ $:=\left\{\operatorname{GSN}_{kl}^{S}\geq-2\sqrt{\frac{2\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\},$ (56)
@@
Here, we use the separation condition to bound the term. If we condition on $\overline{\mathcal{E}}_{S,l}^{\ast}$ and $\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}$, then $\mathrm{IP}_{kli}^{S}$ has a univariate Gaussian distribution with mean zero and variance as the following:
@@
$\overset{(i)}{=}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+\underbrace{\left\|\overline{\mathcal{E}}_{S,l}^{\ast}-(1-|G_{k}^{\ast}|^{-1})^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right\|_{2}^{2}}_{:=\operatorname{CN}_{kli}^{S}}$
@@
$\mathcal{B}^{(i,5)}_{kl,S}$ $:=\left\{\operatorname{CN}_{kli}^{S}\leq(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})\right\},$
$\mathcal{B}^{(i,6)}_{kl,S}$ $:=\left\{\operatorname{GN}_{kli}^{S}\leq 2\sqrt{2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})\zeta^{S}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\}.$
@@
$D^{S}_{kli}(\mathbf{X})$$\gtrsim$ $~\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}+\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl(|S|-2\sqrt{|S|\zeta^{S}}\bigr)$
@@
$\gtrsim$ $~\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}-\sqrt{\frac{\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}$
@@
$~=$ $~\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}~-~\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)|S|-R^{S}_{kl},$
@@
uniformly over $S\in\mathcal{S}$, with probability at least 1C/n1-C/n, where $C>0$ is a constant. This completes the proof of Lemma 19.
@@ C.4.6 Proof of Lemma 20
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-13.00005pt\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}\leq\sup_{\mathbf{v}\in\mathbb{R}^{n},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}=\|\mathcal{E}_{S,\cdot}\|_{\mathrm{op}}^{2}.$
@@ C.4.7 Proof of Lemma 21
$\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\hskip-4.49997ptD^{S}_{kli}(\mathbf{X})\gtrsim\underbrace{\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\hskip-4.49997pt\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|}_{\text{Part A}}\underbrace{-\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\hskip-4.49997ptR^{S}_{kl}}_{\text{Part B}},$
@@
$\min_{1\leq k\neq l\leq K}\left\{\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\hskip-4.49997ptD^{S}_{kli}(\mathbf{X})\right\}$ $\gtrsim\min_{1\leq k\neq l\leq K}\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\hskip-4.49997pt\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|$
@@ C.4.8 Proof of Lemma 22
$t^{S}_{lk}(\dot{\lambda}^{S})$ $\overset{(i)}{=}\sum_{j\in G_{l}^{\ast}}\biggl\{\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\dot{\lambda}^{S}\biggr\}$
$\overset{(ii)}{=}\sum_{j\in G_{l}^{\ast}}\biggl\{\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\bigl(|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}\bigr)\biggr\}$
@@
where step $(i)$ follows from the definition of $t^{S}_{lk}(\lambda^{S})$ in (30), step $(ii)$ follows from the definition of $\dot{\lambda}_{S}$ in Lemma 21, step $(iii)$ uses $\Delta_{S\cap S_{0}}^{2}>-\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}$, step $(iv)$ uses
$m\geq-2\left\{\frac{|G_{k}^{\ast}|\cdot|G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\},$
@@ C.4.9 Proof of Lemma 23
$\frac{1}{\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}}\bigl\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\mathcal{E}_{S\cap S_{0},i}\bigr\rangle\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1),\quad i=1,\ldots,|G_{k}^{\ast}|,$
@@
$\mathbb{P}\bigl(\bigcup_{1\leq k\neq l\leq K}\bigcup_{S\in\mathcal{S}}(\mathcal{C}_{Ikl}^{S})^{C}\bigr)\leq\hskip-3.99994pt\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{Ikl}^{S})^{C})\overset{(i)}{\leq}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\frac{1}{n|S|^{2}{\binom{p}{|S|}}}=\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\frac{1}{nq^{2}}\leq\frac{\pi^{2}}{6n},$
@@ C.4.11 Proof of Lemma 25
$G^{S}_{1,kl}(\mathbf{v})$ $=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-10.00002ptv_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$ (63)
$=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-10.00002pt\frac{1}{2}(v_{i}+v_{j})\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$
@@
$\mathbb{E}\bigl[\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|},\mathbf{X})\bigr]=\mathbb{E}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-10.00002ptv_{i}\mathbb{E}\bigl[\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\bigr]=\sum_{i\in G_{k}^{\ast}}v_{i}\mathbb{E}\bigl[\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr]\overset{(i)}{=}|S|\sum_{i\in G_{k}^{\ast}}v_{i}=0,$
@@
We start by bounding $\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}$. First note that for $\mathbf{M}\in\mathcal{A}^{S}_{kl}$, its operator norm can be simplified into:
@@
$=\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-10.00002ptu_{i}u_{j}\frac{v_{i}+v_{j}}{2}$
@@
$\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}},\mathbf{X})=\hskip-10.00002pt\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}\hskip-10.00002ptv_{j}\bigl\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr\rangle+\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}.$
@@
$\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}\hskip-10.00002ptv_{j}\bigl\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr\rangle$ $=\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl|\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}\hskip-10.00002ptv_{j}\bigl\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr\rangle\bigr|$
@@
$\mathbb{E}\biggl[\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}\hskip-10.00002ptv_{j}\bigl\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr\rangle\biggr]$ $\leq\mathbb{E}\biggl[\bigl(\sum_{j\in G_{k}^{\ast}}\bigl\langle\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr\rangle^{2}\bigr)^{1/2}\biggr]$
@@
$\stackrel{{\scriptstyle(ii)}}{{=}}\bigl(\sum_{j\in G_{k}^{\ast}}\mathbb{E}\|\thinspace\sum_{i\in G_{k}^{\ast},i\neq j}\thinspace\mathcal{E}_{S,i}\|_{2}^{2}\;\bigr)^{1/2}$
@@
$\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl|\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr|$ $\overset{(i)}{=}\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl|\sum_{i\in G_{k}^{*}}v_{i}\bigr(\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\bigl)\bigr|$
@@
$\mathcal{C}_{N,kl}^{S}:=\biggl\{\sup_{\mathbf{v}\in\mathbb{V}_{k}}\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-10.00002ptv_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\lesssim\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\underbrace{|G_{k}^{\ast}|\sqrt{|S|}}_{\hskip-81.94473pt\mathbb{E}\bigl[\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\bigr]\lesssim}~+~\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\zeta^{S}_{N,kl}\biggr\},$
@@
$\mathbb{P}\biggl((\mathcal{C}_{N,kl}^{S})^{C}\biggr)$ $\lesssim\exp\biggl[-\min\biggl(\underbrace{\frac{(\zeta^{S}_{N,kl})^{2}}{|G_{k}^{\ast}|^{2}|S|}}_{\hskip-57.36131pt\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{\mathcal{A}^{S}_{kl}}^{2}\lesssim},\quad\underbrace{\frac{\zeta^{S}_{N,kl}}{\sqrt{|G_{k}^{\ast}|}}}_{\hskip-24.58342pt\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}\lesssim}\biggr)\biggr]$